# Trilinos: Enabling Scientific Computing Across Diverse Hardware Architectures at Scale

### v1

@@ 5.2. Local Field Evaluation: Phalanx
Phalanx is a local field evaluation library designed for equation assembly in PDE applications. The goal of Phalanx is to decompose a complex problem into a number of simpler problems with managed dependencies to support rapid development and extensibility of PDE codes (Notz et al., 2012; Pawlowski et al., 2012a, b). The data structures use Kokkos for performance portability. Through the use of template metaprogramming, Phalanx supports arbitrary user defined data types and evaluation types. This feature allows for simple integration of automatic differentiation tools via operator-overloaded scalar types from the Sacado package (Section 4.3). From a simple definition of equations, quantities such as Jacobians, Jacobian-vector products, Hessians, and parameter sensitivities can be evaluated to machine precision. These quantities can be used by other Trilinos packages for operations including Newton-based nonlinear solves, gradient-based optimization, constraint enforcement, and bifurcation analysis (Pawlowski et al., 2012a, b).
@@ 5.3. Time Integration: Tempus
    • Classic one-step methods (e.g., forward Euler and trapezoidal methods)
    • Explicit Runge-Kutta (RK) methods (e.g., RK explicit 4 Stage)
    • Diagonally Implicit Runge-Kutta (DIRK) methods (e.g., general tableau DIRK and many specific DIRK/SDIRK methods)
    • Implicit-Explicit (IMEX) Runge-Kutta methods (e.g., IMEX RK SSP2, IMEX RK SSP3, and general tableau IMEX RK methods)
    • Multi-Step methods (i.e., BDF2)
    • Second-order ODE methods (e.g., Leapfrog, Newmark methods, and HHT-Alpha)
    • Steppers with subSteppers (e.g., operator-split and subcycling methods)
@@ 5.4. Finite Element Analysis: Panzer
Panzer can provide low level utilities for application codes to build on, or can be used as a high-level application framework. Important capabilities include the following:
@@ 6.1. Build and Test Infrastructure
The Trilinos Framework infrastructure is built on top of the Tribal Build, Integration, and Test System (TriBITS (Bartlett, 2014)), which is built on top of the open-source tools CMake and CTest [^11] https://cmake.org. The TriBITS framework allows building arbitrary subgraphs of dependent (Trilinos) CMake packages in one or more individual aggregated CMake projects (in any arrangement desired). Each Trilinos/TriBITS package lists its direct (required and optional) dependent upstream packages, thus forming a package dependency graph. The TriBITS framework uses this package dependency graph to automatically determine what indirect dependent internal packages must be enabled and processed (and built) and what external packages must be found. TriBITS then orchestrates the processing of all of the required CMake code to find the needed external packages and configure and build (and optionally test and install) the selected set of internal packages. This allows a large number of (Trilinos) CMake packages to be configured, built, and tested in a flexible and efficient manner. In addition, TriBITS provides support for a number of advanced features that are not available in raw CMake/CTest including: eliminating a large amount of boiler-plate CMake code and avoiding common mistakes; enabling and testing all downstream packages given a set of enabled (i.e. modified) upstream packages; managing the enabling and disabling of tests based on various criteria; producing build and test results submitted to a CDash site on a package-by-package basis; producing reduced source tarballs for only a desired subset of enabled packages. As of version 14.4 of Trilinos, TriBITS and Trilinos have been updated to allow packages using raw CMake to be integrated with just a few well-defined integration requirements. The TriBITS framework has allowed Trilinos to scalably grow in the number of packages and the complexity without undue burden on individual Trilinos developers and users.
@@ 7.1. Contributing to Trilinos
Contributions to Trilinos can be offered through the standard GitHub pull request model. Proposed code changes are required to pass a set of tests as well as review and approval prior to be merged. Detailed instructions can be found in the contributing guidelines [^15] https://github.com/trilinos/Trilinos/blob/master/CONTRIBUTING.md in the source code repository.
@@ 7.2. Platforms for exchange among users or developers
The community of Trilinos users and developers operates several forums for exchange and discussion. Technical discussions about the source code and its development happen within the Trilinos GitHub repository [^16] https://github.com/trilinos/Trilinos. Several mailing lists (see https://trilinos.github.io/mail_lists.html) distribute relevant information and updates on Trilinos. The #trilinos channel within the Kokkos slack workspace [^17] See https://kokkos.org for details. provides a quick and accessible forum to ask questions.
@@ 8. Concluding remarks
The continued existence and on-going development of the Trilinos project demonstrates its critical role in scientific and high-performance computing. This update underscores Trilinos’ dedication to maintaining relevance amidst the rapid development of HPC software frameworks and hardware architectures, and the demands of increasingly complicated multiscale multiphysics simulation code bases for tackling scientific and engineering problems. By adopting the Kokkos ecosystem, expanding package functionalities, and embracing a modular structure, Trilinos ensures both performance portability and adaptability to existing and emerging application domains.

### v2

@@ 5.2. Local Field Evaluation: Phalanx
Phalanx is a local field evaluation library designed for equation assembly in PDE applications. The goal of Phalanx is to decompose a complex problem into a number of simpler problems with managed dependencies to support rapid development and extensibility of PDE codes (Notz et al., 2012; Pawlowski et al., 2012b, a). The data structures use Kokkos for performance portability. Through the use of template metaprogramming, Phalanx supports arbitrary user defined data types and evaluation types. This feature allows for simple integration of automatic differentiation tools via operator-overloaded scalar types from the Sacado package (Section 4.3). From a simple definition of equations, quantities such as Jacobians, Jacobian-vector products, Hessians, and parameter sensitivities can be evaluated to machine precision. These quantities can be used by other Trilinos packages for operations including Newton-based nonlinear solves, gradient-based optimization, constraint enforcement, and bifurcation analysis (Pawlowski et al., 2012b, a).
@@ 5.3. Time Integration: Tempus
    • Classic one-step methods (e.g., forward Euler and trapezoidal methods),
    • Explicit Runge-Kutta (RK) methods (e.g., RK explicit 4 Stage),
    • Diagonally Implicit Runge-Kutta (DIRK) methods (e.g., general tableau DIRK and many specific DIRK/SDIRK methods),
    • Implicit-Explicit (IMEX) Runge-Kutta methods (e.g., IMEX RK SSP2, IMEX RK SSP3, and general tableau IMEX RK methods),
    • Multi-Step methods (i.e., BDF2),
    • Second-order ODE methods (e.g., Leapfrog, Newmark methods, and HHT-Alpha),
    • Steppers with subSteppers (e.g., operator-split and subcycling methods).
@@ 5.4. Finite Element Analysis: Panzer
Panzer can provide utilities for application codes to build on, or can be used as a high-level application framework. Important capabilities include the following.
@@ 6.1. Build and Test Infrastructure
The Trilinos Framework infrastructure is built on top of the Tribal Build, Integration, and Test System (TriBITS (Bartlett, 2014)), which is built on top of the open-source tools CMake and CTest [^14] https://cmake.org. The TriBITS framework allows building arbitrary subgraphs of dependent (Trilinos) CMake packages in one or more individual aggregated CMake projects (in any arrangement desired). Each Trilinos/TriBITS package lists its direct (required and optional) dependent upstream packages, thus forming a package dependency graph. The TriBITS framework uses this package dependency graph to automatically determine what indirect dependent internal packages must be enabled and processed (and built) and what external packages must be found. TriBITS then orchestrates the processing of all of the required CMake code to find the needed external packages and configure and build (and optionally test and install) the selected set of internal packages. This allows a large number of (Trilinos) CMake packages to be configured, built, and tested in a flexible and efficient manner. In addition, TriBITS provides support for a number of advanced features that are not available in raw CMake/CTest including: eliminating a large amount of boiler-plate CMake code and avoiding common mistakes; enabling and testing all downstream packages given a set of enabled (i.e. modified) upstream packages; managing the enabling and disabling of tests based on various criteria; producing build and test results submitted to a CDash site on a package-by-package basis; producing reduced source tarballs for only a desired subset of enabled packages. As of version 14.4 of Trilinos, TriBITS and Trilinos have been updated to allow packages using raw CMake to be integrated with just a few well-defined integration requirements. The TriBITS framework has allowed Trilinos to scalably grow in the number of packages and the complexity without undue burden on individual Trilinos developers and users. An always up-to-date overview of the configurations used for automated testing of pull requests against Trilinos can be found in the source code repository [^15] https://github.com/trilinos/Trilinos/blob/master/.github/workflows/AT2.yml.
@@ 7.1. Contributing to Trilinos
Contributions to Trilinos can be offered through the standard GitHub pull request model. Proposed code changes are required to pass a set of tests as well as review and approval prior to be merged. Detailed instructions can be found in the contributing guidelines [^19] https://github.com/trilinos/Trilinos/blob/master/CONTRIBUTING.md in the source code repository. Questions regarding the inclusion of new packages should be discussed with the Trilinos leadership team [^20] https://trilinos.github.io/team.html.
@@ 7.2. Platforms for exchange among users or developers
The community of Trilinos users and developers operates several forums for exchange and discussion. Technical discussions about the source code and its development happen within the Trilinos GitHub repository [^21] https://github.com/trilinos/Trilinos. The #trilinos channel within the Kokkos slack workspace [^22] See https://kokkos.org for details. provides a quick and accessible forum to ask questions.
@@ 8. Concluding remarks
The evolution of the Trilinos project demonstrates its critical role in scientific and high-performance computing. This update underscores Trilinos’ dedication to maintaining relevance amidst the rapid development of HPC software frameworks and hardware architectures, and the demands of increasingly complicated multiscale multiphysics simulation code bases for tackling scientific and engineering problems. By adopting the Kokkos ecosystem, expanding package functionalities, and embracing a modular structure, Trilinos ensures both performance portability and adaptability to existing and emerging application domains.